# Metis-SPECS: Decoupling Learning via Self-distilled Preference-based Cold Start for VLMs

### v1

@@ 4.4 Analysis of the relationship between GF and Final Performance
4.5 Analysis of the impact on RL Training Efficiency and Stability
Performance and Training Efficiency. To evaluate performance and training efficiency, we tracked MEGA-Bench scores throughout the GRPO training process. As illustrated in Figure 4, the DPO+GRPO model begins with a substantially higher initial score, demonstrating the immediate benefit of preference-based pre-alignment. Furthermore, it maintains a clear advantage throughout training, converging more rapidly and ultimately achieving a higher performance ceiling than its SFT+GRPO counterpart. For more analysis on this content, please refer to Appendix A.3.
Training Stability. Beyond performance metrics, we analyzed training stability by comparing the policy loss curves, presented in Figure 4. The curve for DPO+GRPO is visibly smoother and more stable, indicating a more consistent and reliable optimization trajectory. In contrast, the SFT+GRPO policy exhibits greater volatility, suggesting that the RL algorithm must make more drastic and potentially erratic updates. In terms of format rewards, RL based on SFT cold start is also weaker than RL based on DPO cold start in terms of the stability of format rewards.
@@ 5 Related Work
For a more detailed introduction to related work, please refer to Appendix A.5.
RL for Multimodal Reasoning. The application of RL has proven to be a highly effective method for enhancing the reasoning capabilities of large language models, with notable successes such as DeepSeek-R1 (Guo et al., 2025) which leverages RLVR (Lambert et al., 2024; Guo et al., 2025). Inspired by these advancements in the text-only domain, a substantial and rapidly growing body of research has begun to adapt these RL techniques for VLMs. This has catalyzed a wave of ”MLLM-r1” studies (Meng et al., 2025; Shen et al., 2025; Peng et al., 2025; Zhou et al., 2025; Zhang et al., 2025a; Wang et al., 2025b; a; Zheng et al., 2025; Ma et al., 2025), all aiming to harness similar principles to unlock more advanced multimodal reasoning abilities.
Cold-Start Strategies and Data Generation. A crucial precursor to effective reinforcement learning is the ”cold-start” phase, which initializes the model policy. The predominant strategy for this phase is SFT on high-quality instruction-response pairs (Wei et al., 2025; Yang et al., 2025; Huang et al., 2025; Deng et al., 2025b). In parallel, the prohibitive cost of human annotation has driven the field towards synthetic data generation, often using powerful teacher models to distill data for smaller student models (Zhang et al., 2025b; Yao et al., 2024; Xu et al., 2024; Huang et al., 2025).
@@ 6 Conclusions
To ensure the reproducibility of our work, we provide a detailed account of our methodology and experimental setup. The core SPECS framework, including its three stages of self-distillation for data generation, DPO-based pre-alignment, and final GRPO fine-tuning, is described in Section 3. Our complete experimental settings, including the datasets, benchmarks, and baselines used for evaluation, are detailed in Section 4.1 . This section also specifies crucial implementation details, such as learning rates and batch sizes for all training stages . We utilized publicly available frameworks, MM-EUREKA and LlamaFactory, for our implementation. Furthermore, the appendix offers additional resources to aid in reproduction, including the exact system prompts used for training and inference (Appendix A.2) and a detailed analysis of the hybrid loss coefficient (Appendix A.3).
Appendix A Appendix
A.1 The Use of Large Language Models
In preparing this manuscript, we utilized the Large Language Model (LLM) Gemini-2.5-pro (Comanici et al., 2025) to polishing the text. Its application was strictly limited to correcting spelling and grammatical errors. The authors manually reviewed and verified all AI-assisted modifications to ensure factual accuracy. The core ideas, methodologies, and figures presented are entirely the original work of the human authors.
A.2 Prompts
@@
System Prompt for Cold Start Training and ID Task Inference The following is the Inference System Prompt for the OOD Task in the instruction format generalization experiment of Section 2. System Prompt for OOD Task Inference
A.3 Analysis of Loss Balance Coefficients in Cold-Start Training
A.3.1 Preliminaries
@@
Direct Preference Optimization. DPO(Rafailov et al., 2023) refines models by learning directly from preference data, consisting of a prompt $x$, a chosen response $y_{c}$, and a rejected response $y_{r}$. The DPO loss function is designed to increase the relative probability of the chosen response over the rejected one: L_DPO(π_θ; π_ref) = -E_(x, y_c, y_r) ∼D [ logσ( βlogπθ(yc—x)πref(yc—x) - βlogπθ(yr—x)πref(yr—x) ) ]
@@
A.3.2 Analysis of Loss Balance Coefficients
@@
Among the $\mathcal{L}_{hybrid}$, the DPO loss function is designed to increase the relative probability of the chosen response over the rejected one: L_DPO(π_θ; π_ref) = -E_(x, y_c, y_r) ∼D [ logσ( βlogπθ(yc—x)πref(yc—x) - βlogπθ(yr—x)πref(yr—x) ) ]
@@
A.4 Performance Potential of DPO-based Cold-Start vs. SFT
@@
Exploration Capability via Rollout Branching Factor (RBF). Beyond task performance, we investigated the intrinsic exploratory capacity of models, a critical factor for successful reinforcement learning. We measure this using the RBF (Deng et al., 2025a), which quantifies the diversity of a model’s generation by counting the number of candidate tokens within the probability mass of top p (0.95) during decoding. A higher RBF signifies greater generation diversity and, consequently, a stronger capacity for exploration. As shown in Table 4, our DPO-based cold-start method yields a substantially higher RBF than the SFT baseline. This finding suggests that our approach cultivates a broader exploration space, which is highly beneficial for the subsequent RL phase, enabling the model to discover more diverse and potentially higher-quality solutions. This outcome highlights a key advantage of our method in preparing models for alignment.
A.5 Related Work
The application of RL has emerged as a highly effective method for enhancing the reasoning capabilities of large language models, with notable successes in the text-only domain such as DeepSeek-R1 (Guo et al., 2025), which leverages RLVR (Lambert et al., 2024; Guo et al., 2025). Inspired by these advancements, a substantial and rapidly growing body of research has begun adapting RL techniques for Vision-Language Models (VLMs). This has catalyzed a wave of ”MLLM-r1” studies, all aiming to harness similar principles to unlock more advanced multimodal reasoning abilities. For instance, specific approaches include exploring the frontiers of multimodal reasoning with rule-based reinforcement learning in MM-Eureka (Meng et al., 2025), creating a stable and generalizable R1-style model with VLM-R1 (Shen et al., 2025), and empowering smaller models using two-stage rule-based RL as seen in LMM-R1 (Peng et al., 2025). Researchers have also demonstrated significant visual reasoning on non-SFT models with R1-Zero (Zhou et al., 2025), learned step-wise reasoning with R1-VL (Zhang et al., 2025a), and achieved data-efficient self-improvement using MCTS-guided sample selection in SOTA with less (Wang et al., 2025b). Furthermore, innovative methods seek to incentivize self-reflection with VL-Rethinker (Wang et al., 2025a), encourage ”thinking with images” in DeepEyes (Zheng et al., 2025), or unify various visual tasks under a single RL framework as proposed in One RL to See Them All (Ma et al., 2025).
A crucial precursor to effective reinforcement learning is the “cold-start” phase, which initializes the model’s policy before the RL stage begins. The predominant strategy for this phase is SFT on high-quality instruction-response pairs, a foundational step adopted by many leading models to establish a strong baseline performance (Wei et al., 2025; Yang et al., 2025; Huang et al., 2025; Deng et al., 2025b). In parallel with refining cold-start methods, the prohibitive cost of human annotation has driven the field towards synthetic data generation. This approach often involves using powerful teacher models to distill vast amounts of data for training smaller student models. This trend is evident in a variety of techniques, such as synthesizing multimodal instruction data from a single image with Oasis (Zhang et al., 2025b), empowering models with reflection via collective Monte Carlo Tree Search in Mulberry (Yao et al., 2024), and enabling models to reason step-by-step with LLaVA-CoT (Xu et al., 2024). These data-centric strategies, including those used to generate data for models like Vision-R1 (Huang et al., 2025), are vital for creating the rich, high-quality datasets needed to effectively train and align modern VLMs for complex reasoning tasks.
A.6 Qualitative Examples

### v2

@@ 5 Related Work
The application of RL has emerged as a highly effective method for enhancing the reasoning capabilities of large language models, with notable successes in the text-only domain such as DeepSeek-R1 (guo2025deepseek), which leverages RLVR (lambert2024tulu; guo2025deepseek). Inspired by these advancements, a substantial and rapidly growing body of research has begun adapting RL techniques for VLMs. This has catalyzed a wave of “MLLM-r1” studies, all aiming to harness similar principles to unlock more advanced multimodal reasoning abilities. For instance, MM-Eureka (meng2025mmeureka) explores the enhancement of multimodal reasoning abilities through rule-based RL by constructing high-quality multimodal reasoning datasets. VL-Rethinker (wang2025vl) stimulates the slow thinking and self-reflection abilities of VLMs through RL. Orsta (ma2025one) establishes a unified RL system that supports VLMs in jointly learning visual reasoning and perception tasks. VLM-R1 (shen2025vlm) extends R1-style RL to VLMs for visual understanding tasks to improve their visual reasoning abilities. LMM-R1 (peng2025lmm) enhances the model’s basic reasoning ability and multimodal generalization ability through a two-stage training strategy of basic reasoning enhancement and multimodal generalization training. R1-VL (zhang2025r1) realizes the self-improvement of MLLMs’ reasoning ability by solving the sparse reward problem through Step-wise GRPO. DeepEyes (zheng2025deepeyes) motivates the model’s “Thinking with Images” ability through RL. VisualThinker-R1-Zero (zhou2025r1) performs RL directly without any supervised fine-tuning of the model to reproduce the “aha moment”.
A crucial precursor to effective RL is the “cold-start” phase, which initializes the model’s policy before the RL stage begins. The conventional strategy for this phase is SFT, a foundational step adopted by many leading models to establish a strong baseline performance (wei2025advancing; yang2025r1; huang2025vision; deng2025openvlthinker). In parallel with refining cold-start methods, the prohibitive cost of human annotation has driven the field towards synthetic data generation. This approach often involves using powerful teacher models to distill vast amounts of data for training smaller student models (zhang2025oasis; xu2024llava; huang2025vision). Vision-R1 (huang2025vision) cold-starts the model before applying RL by synthesizing 100K high-quality long CoT instructions. LLaVA-CoT (xu2024llava) integrates multiple mainstream visual question answering datasets and uses advanced large models to synthesize 99K valid image-question-answer pairs. R1-Onevision (yang2025r1) adopts a two-stage training strategy of SFT + RL by synthesizing a 155K instruction set.
@@ 6 Conclusions
To ensure the reproducibility of our work, we provide a detailed account of our methodology and experimental setup. The core SPECS framework, including its three stages of self-distillation for data generation, DPO-based pre-alignment, and final GRPO fine-tuning, is described in Section 3. Our complete experimental settings, including the datasets, benchmarks, and baselines used for evaluation, are detailed in Section 4.1 . This section also specifies crucial implementation details, such as learning rates and batch sizes for all training stages . We utilized publicly available frameworks, MM-EUREKA and LlamaFactory, for our implementation. Furthermore, the appendix offers additional resources to aid in reproduction, including the exact system prompts used for training and inference (Appendix B) and a detailed analysis of the hybrid loss coefficient (Appendix C).
Appendix A The Use of Large Language Models
In preparing this manuscript, we utilized the Large Language Model (LLM) Gemini-2.5-pro (comanici2025gemini) to polishing the text. Its application was strictly limited to correcting spelling and grammatical errors. The authors manually reviewed and verified all AI-assisted modifications to ensure factual accuracy. The core ideas, methodologies, and figures presented are entirely the original work of the human authors.
Appendix B Prompts
@@
System Prompt for Cold Start Training and ID Task Inference The following is the Inference System Prompt for the OOD Task in the instruction format generalization experiment of Section 2. System Prompt for OOD Task Inference The following Prompt is used for filtering chosen responses, see Section 3.2. System Prompt for Chosen Response Filtration
Appendix C Analysis of Loss Balance Coefficients in Cold-Start Training
C.1 Preliminaries
@@
Direct Preference Optimization. DPO(rafailov2023direct) refines models by learning directly from preference data, consisting of a prompt $x$, a chosen response $y_{c}$, and a rejected response $y_{r}$. The DPO loss function is designed to increase the relative probability of the chosen response over the rejected one: L_DPO(π_θ; π_ref) = -E_(x, y_c, y_r) ∼D [ logσ( βlogπθ(yc—x)πref(yc—x) - βlogπθ(yr—x)πref(yr—x) ) ]
@@
C.2 Analysis of Loss Balance Coefficients
@@
Among the $\mathcal{L}_{hybrid}$, the DPO loss function is designed to increase the relative probability of the chosen response over the rejected one: L_hybrid = -E_(x, y_c, y_r) ∼D [ logσ( βlogπθ(yc—x)πref(yc—x) - βlogπθ(yr—x)πref(yr—x) ) ] + λE_(x, y_c) ∼D [-logπ_θ(y_c—x)]
@@
Appendix D Performance Potential of DPO-based Cold-Start vs. SFT
@@
Exploration Capability via Rollout Branching Factor (RBF). Beyond task performance, we investigated the intrinsic exploratory capacity of models, a critical factor for successful RL. We measure this using the RBF (deng2025trial), which quantifies the diversity of a model’s generation by counting the number of candidate tokens within the probability mass of top p (0.95) during decoding. A higher RBF signifies greater generation diversity and, consequently, a stronger capacity for exploration. As shown in Table 6, our DPO-based cold-start method yields a substantially higher RBF than the SFT baseline. This finding suggests that our approach cultivates a broader exploration space, which is highly beneficial for the subsequent RL phase, enabling the model to discover more diverse and potentially higher-quality solutions. This outcome highlights a key advantage of our method in preparing models for alignment.
Appendix E Analysis of Chosen Response Filtration
We believe it is crucial to ensure the quality and accuracy of chosen response in cold start parse. However, we found that there are a small number of cases (generated by GRPO-zero) in chosen response where the answers are not consistent with the reasoning, which could affect training, such as the following case:
[…]
The reasons for this situation still need to be explored, and we think it may be related to reward hacking in GRPO training (because the answer to this training data was incorrectly labeled as 3, while the correct answer should be 4 as in the reasoning process). To prevent these data from affecting the cold start training, we use another LLM to verify the data. This evaluation process is simple and objective, with no significant errors, and will not introduce bias into data filtering. The table 7 below shows the differences in the final model performance between the data filtered by the advanced LLM and the randomly filtered data.
Table 7: Performance Comparison of Self-Distillation with and without Verification on Various Benchmarks
Model Megabench MMMU MathVista MathVision MathVerse AVG
Qwen2.5-7B-Instruct 35.07 54.20 63.70 25.40 38.20 43.31
w.o verification 38.72 55.11 73.50 26.90 46.44 48.13
w verification 39.17 56.78 75.90 29.50 48.73 50.02
It can be seen from the experimental results that the model after data filtering performs better than the final model obtained by randomly selecting data. On the other hand, regarding the selection of large models for screening, in addition to using some advanced large models, due to the simplicity of the task instructions, some open-source large models can also be used as alternatives.
Appendix F Experiments Implementation Details
All experiments were conducted on a cluster of 32 NVIDIA H800 (80G) GPUs. We utilize the open-source Multimodal Large Language Model, Qwen2.5-VL-7B (Qwen2.5-VL), as our base model.
For the GRPO training in Stage 1 and Stage 3, we employ the MM-EUREKA  [^7] https://github.com/ModalMinds/MM-EUREKA framework. The rollout and training batch sizes are both set to $128$, with $8$ rollouts generated per sample. The learning rate is configured to $1\times 10^{-6}$. Following the DAPO (yu2025dapo) methodology, we set the clipping thresholds to $0.2$ (lower) and $0.28$ (upper), and we do not apply a KL penalty. The maximum output length is restricted to $10,240$ tokens. This training phase spanned $400$ steps and was completed in approximately $12$ hours.
For the DPO training in Stage 2, as well as for the comparative SFT experiments, we leverage the LlamaFactory  [^8] https://github.com/hiyouga/LLaMA-Factory framework using a total dataset size of $9,155$ samples. In this configuration, the training batch size is set to $64$, the learning rate is maintained at $1\times 10^{-6}$, and the hyperparameter $\lambda$ for the hybrid loss function is set to $1$. The maximum output length is extended to $16,384$ tokens. These models were trained for $140$ steps, with the training process completing in under $30$ minutes. The prompt used during training is shown in Appendix B.
We have counted the computing time for DPO and SFT training using 8 GPUs under the same training parameters (including steps, batch size, data length, max new tokens, etc.), as shown in the following table 8.
Table 8: Comparison of training performance between SFT and DPO.
Training Method Train Samples/s Train Steps/s Total FLOS Run Time/s
SFT 1.035 0.129 $2.82\times 10^{17}$ 8848
DPO 0.979 0.123 $6.22\times 10^{17}$ 9346
$\Delta$ -0.056 -0.006 $3.4\times 10^{17} +$498
As can be seen from the table, in actual computation, the computational cost between SFT and DPO is not significantly different. For the computation of 9,000 pieces of data, the total computation time only increased by approximately 8.3 minutes.
Appendix G Qualitative Examples